# Lumos-1: On Autoregressive Video Generation from a Unified Model Perspective

### v1

@@ 4.3 Analysis and Ablation Studies
Qualitative visual comparison. We compare Lumos-1 with popular video generation methods in figure 6. For T2V, the visual quality of our 384p videos does not trail 512p videos from LTX-Video. In the provided case, Lumos-1 generates better natural motion (water waves) and aligns with prompts better (skier in red and waves). For I2V, Lumos-1 handles multiple-object (multiple floating hot air balloons in example 1) and fine-grained motion (subtle ripples around the shoreline in example 3) substantially better than Stable Video Diffusion [5] (SVD), which generates global camera movement only. In example 2, SVD produces significant blurs, whereas Lumos-1 animates subjects smoothly. More visualizations are placed in the Appendix.
Effectiveness of temporal-tube masks during AR-DF training. In figure 7(a), we compare the frame-wise validation loss (frame 0, 3, 6) when using global random masks and temporal tube masks. The training mask ratios $\rho_{tra}$ are both set to $0.7$ following MAR [33]. For random masks, the loss of frame 6 decreases immediately and becomes lower than that of earlier frames. This rapid fall indicates pronounced information leakage: the model can reconstruct a masked token by attending to unmasked tokens in neighbouring frames instead of modelling genuine temporal dynamics, rendering the task overly easy. For temporal tube masks, frame 6 is the hardest because all pixels in the same spatial locations are masked across the temporal axis, eliminating the shortcut available with random masks. As iterations proceed, the gap between frames narrows and eventually levels out, demonstrating that the model is learning to propagate information through time rather than copying it.
Figure 7: (a) Effectiveness of temporal-tube masks during AR-DF training (on videos) using the 0.5B model; (b) Effectiveness of MM-RoPE; (c) Sensitivity analysis of CFG scale on GenEval using 1B model.
@@
Effect of AR-DF inference masks and sensitivity to $\rho_{inf}$. AR-DF requires the same partial-context masking at inference as during training; omitting these masks severely harms quality. In figure 8(b), we observe that the “w/o inference mask" setting produces visible artifacts and flickering, whereas using masks preserves coherence. In figure 8(a), we select two metrics (imaging quality and dynamic degree) from VBench to quantitatively assess the impact of $\rho_{inf}$. A broad plateau between $0.3$ and $0.7$ yields smooth, visually pleasing videos. When $\rho_{inf}$ is below $0.3$, insufficient context forces the model to degrade both motion and per-frame quality, thus inflating dynamic degree values. When $\rho_{inf}$ is above $0.7$, overly aggressive masking disrupts temporal continuity, driving up dynamic degree values. We empirically set $\rho_{inf}$ to $0.7$ to ensure obvious motion.
Table 4: Inference time with different RoPEs using 1B / 3B model. Results are reported on one H20 with batch size set to 1, $N_{steps}=50$ and CFG used by default.
1D RoPE M-RoPE MM-RoPE
@@
Effectiveness of MM-RoPE. figure 7(b) plots the validation loss of the 0.5B model under four RoPE settings. Note that M-RoPE means that both designs are removed. We can observe that MM-RoPE consistently converges faster and settles at the lowest loss, confirming the benefit of modelling fine-grained spatiotemporal information. Although ablating either component raises the loss, dropping the distributed design hurts more than dropping the design of scaled position, indicating that comprehensive frequency allocation is the dominant factor. Removing both enhancements gives the slowest convergence and the highest plateau, showing that the two mechanisms are complementary for effective video generation.
Figure 9: (a) Effect of the number of meta MM-RoPEs in MM-RoPE; (b) Effect of the scaling factors in MM-RoPE. Both curves are plotted using the 0.5B model trained on videos.
Effect of the number of meta MM-RoPEs in MM-RoPE. MM-RoPE slices the embedding channels into several meta groups. More groups mean that one kind of information (temporal, height or width) receives a wider and more comprehensive frequency spectrum for modeling. figure 9(a) plots the validation loss for a 0.5B model under four settings: 1) w/o distributed design: This setting uses the previous design, which allocates the first $\frac{2}{8}$ channels for temporal modeling and $\frac{3}{8}$ channels for height and width modeling, respectively. 2) #meta MM-RoPE=1: This setting equips a meta MM-RoPE of 64 channels, while maintaining the ratios ($2:3:3$) for modeling temporal, height and width information. This variant improves upon the previous one by enabling interleaved height and width channels, which increases the frequency spectrum range for the two spatial dimensions. 3) #meta MM-RoPE=2: This setting equips a meta MM-RoPE of 32 channels. This variant improves the frequency spectrum range for temporal, height and width information compared with the previous one. 4) #meta MM-RoPE=4: This setting is the default design that keeps the number of channels for every meta MM-RoPE minimal (16 channels). This means that the frequency spectrum range is most comprehensive for the temporal, height, or width dimension. These results confirm that widening the frequency spectrum for every dimension through increasing the number of meta MM-RoPEs substantially improves spatiotemporal modeling and overall training effectiveness.
Effect of the scaling factors in MM-RoPE. figure 9(b) presents the validation loss curves of varying the scaling factors for modeling temporal, height and width positional information in MM-RoPE. Two clear trends emerge: Moving from $(1,1,1)$$(2,4,4)$$(4,8,8)$ steadily lowers the curve, but further enlarging to $(8,16,16)$ or $(16,32,32)$ yields no additional gain since the three curves almost overlap throughout training. Therefore, a moderate scale of $(4,8,8)$ is sufficient to balance vision-language ranges and fully realize the benefit of higher-resolution RoPE, while avoiding unnecessary frequency inflation. We therefore adopt $(4,8,8)$ as the default scaling for MM-RoPE.
Table 5: GenEval score across different aspect ratios.
GenEval 0.605 0.601 0.569
Inference overhead analysis for MM-RoPE. Similar to M-RoPE, MM-RoPE needs to locate the start of visual tokens and then apply the RoPE mechanism, which requires a small amount of computation. In Tab. 4, we compare the inference speed of generating images and videos using the vanilla 1D RoPE, M-RoPE and MM-RoPE. We can observe that: 1) Compared with 1D RoPE, the incorporation of 3D priors introduces only 3.5%-4.1% inference latency. 2) Compared to M-RoPE, MM-RoPE introduces no additional latency.
@@
Sensitivity analysis of CFG scale. We study the impact of the guidance scale on GenEval using 1B model in figure 7(c). We find that scale values from 13 to 16 (default value) lead to decent results.
Robustness to aspect ratios. Although the aspect ratio of training data is mostly $7:4$, Tab. 5 indicates that Lumos-1 1B adapts well to visual generation with different aspect ratios due to the unified codebook design.
@@ 5 Conclusion
In this paper, we introduce Lumos-1, a model that utilizes the LLM architecture for autoregressive video generation. We propose MM-RoPE for better spatiotemporal dynamics modeling and propose AR-DF for effective training and inference considering intra-frame bidirectionality and inter-frame temporal causality. We anticipate that Lumos-1 represents a significant step toward building a foundational unified model.
Acknowledgement
The LaTeX template is built upon Meta’s original template.
@@ Acknowledgement
  • appendix A presents a detailed version of the AR-DF inference algorithm.
  • appendix B presents more details concerning architectures and implementation.
  • appendix C provides more visualizations to demonstrate image and video generation capabilities.
  • appendix D presents discussions covering 1) limitations and future work (Sec. D.1) and 2) potential societal impact and safeguards (Sec. D.2).
Appendix A Detailed AR-DF Inference Algorithm

### v2

@@ 4.3 Analysis and Ablation Studies
Qualitative visual comparison. We compare Lumos-1 with popular video generation methods in figure˜6. For T2V, the visual quality of our 384p videos does not trail 512p videos from LTX-Video. In the provided case, Lumos-1 generates better natural motion (water waves) and aligns with prompts better (skier in red and waves). For I2V, Lumos-1 handles multiple-object (multiple floating hot air balloons in example 1) and fine-grained motion (subtle ripples around the shoreline in example 3) substantially better than Stable Video Diffusion [6] (SVD), which generates global camera movement only. In example 2, SVD produces significant blurs, whereas Lumos-1 animates subjects smoothly. More visualizations are placed in the Appendix.
Effectiveness of temporal‑tube masks during AR‑DF training. In figure˜7(a), we compare the frame-wise validation loss (frame 0, 3, 6) when using global random masks and temporal tube masks. For random masks, the loss of frame 6 decreases immediately and becomes lower than that of earlier frames. This rapid fall indicates pronounced information leakage: the model can reconstruct a masked token by attending to unmasked tokens in neighbouring frames instead of modelling genuine temporal dynamics, rendering the task overly easy. For temporal tube masks, frame 6 is the hardest because it has the longest context to learn from, and all pixels in the same spatial locations are masked across the temporal axis, eliminating the learning shortcut. As iterations proceed, the gap between frames narrows and eventually levels out, demonstrating that the model is learning to propagate information through time rather than copying it.
Figure 7: (a) Effectiveness of temporal-tube masks during AR-DF training (on videos) using the 0.5B model; (b) Effectiveness of $M$-RoPE; (c) Sensitivity analysis of CFG scale on GenEval using 1B model.
@@
Table 6: Inference time analysis with different RoPEs and decoding strategies. Results are reported on one H20 with batch size set to 1, $N_{steps}=50$ and CFG used by default.
1D RoPE (1B / 3B) M‑RoPE (1B / 3B) MM‑RoPE (1B / 3B)
@@
Next-Token Pred Mask Pred w/o KV Cache Mask Pred w/ KV Cache
Video ($448\times 256\times 25$) 960.0s 383.0s 77.8s
Effect of AR‑DF inference masks and sensitivity to $\rho_{inf}$. AR‑DF requires the same partial‑context masking at inference as during training; omitting these masks severely harms quality. In figure˜8(b), we observe that the “w/o inference mask” setting produces visible artifacts and flickering, whereas using masks preserves coherence. In figure˜8(a), we select two VBench metrics (imaging quality and dynamic degree) to assess the impact of $\rho_{inf}$. A broad plateau between $0.3$ and $0.7$ yields smooth, visually pleasing videos. When $\rho_{inf}<0.3$, insufficient context causes motion and per‑frame quality degradation, thus inflating dynamic degree values. When $\rho_{inf}>0.7$, aggressive masking disrupts temporal continuity, driving up dynamic degree values. We set $\rho_{inf}$ to $0.7$ to ensure obvious motion.
Training strategy comparison. Temporal tube masking is a key in AR-DF training. In Tab.˜5, we compare it with other methods, including global random masks, loss reweighting (step decay loss [83] and its modified version, linear decay loss), and diffusion forcing masks, by performing joint training on Lumos-1 1B using 8 H20 for 10k steps. We observe that loss reweighting cannot solve the information leakage issue. Diffusion forcing can ameliorate this, but temporal tube masking in AR-DF is particularly suited for videos, thus surpassing all other methods.
Table 7: Ablation study on $M$-RoPE.
Distributed Scaled 3D GenEval Vbench-OC Vbench-IQ
0.310 0.122 0.350
0.414 0.211 0.451
0.566 0.245 0.535
0.591 0.249 0.559
Scaling Factors GenEval Vbench-OC Vbench-IQ
(1, 1, 1) 0.566 0.245 0.535
(2, 4, 4) 0.583 0.248 0.545
(4, 8, 8) 0.591 0.249 0.559
(8, 16, 16) 0.593 0.250 0.553
@@
Efficacy of MM-RoPE. figure˜7(b) plots the validation loss of the 0.5B model under four RoPE settings. Note that M‑RoPE means that both designs are removed. We can observe that MM‑RoPE consistently converges faster and settles at the lowest loss, confirming the benefit of modelling fine‑grained spatiotemporal information. Although ablating either component raises the loss, dropping the distributed design hurts more than dropping the design of scaled position, indicating that comprehensive frequency allocation is the dominant factor. Removing both enhancements gives the slowest convergence and the highest plateau, showing that the two mechanisms are complementary for effective video generation. We quantitatively ablate on the MM-RoPE design in Tab.˜7 by performing joint training on Lumos-1 1B using 8 H20 for 10k steps. We can find that the distributed design contributes significantly to the boost and speeds up the model’s convergence. By varying the scaling factors, we confirm that (4,8,8) is a suitable and well-justified choice for visual generation.
Inference time analysis. In Tab.˜6, we compare the inference speed using the vanilla 1D RoPE, M-RoPE and MM-RoPE. We can observe that: 1) Compared with 1D RoPE, the incorporation of 3D priors introduces only 3.5%-4.1% inference latency. 2) Compared to M-RoPE, MM-RoPE introduces no additional latency. We also compare the speed using different decoding strategies. We observe a clear efficiency boost compared with vanilla next-token prediction by using discrete diffusion (mask prediction) and KV cache.
Table 8: Performance comparison with other types of RoPE on GenEval and VBench.
RoPE Type GenEval Vbench-OC Vbench-IQ
M-RoPE 0.310 0.122 0.350
U-RoPE 0.402 0.165 0.423
IL-RoPE 0.541 0.225 0.513
VideoRoPE 0.569 0.243 0.540
HoPE 0.570 0.246 0.545
$M$-RoPE 0.591 0.249 0.559
Comparison with different RoPE designs. To validate the efficacy of MM-RoPE, we benchmark it against M-RoPE [80], VideoRoPE [84], U-RoPE [68], IL-RoPE [39] and HoPE [35]. As shown in Tab.˜8, MM-RoPE consistently outperforms them. This superiority stems from its unique design, which holistically incorporates a comprehensive frequency allocation strategy and a strategic scaling for modality alignment and resolution enhancement. A detailed breakdown of each method and a full analysis are provided in Sec.˜E.2.
@@
Sensitivity analysis of CFG scale. We study the impact of the guidance scale on GenEval using 1B model in figure˜7(c). We find that scale values from 13 to 16 (default value) lead to decent results.
@@ 5 Conclusion
In this paper, we introduce Lumos-1, which utilizes the LLM architecture for AR video generation. We propose MM-RoPE for better spatiotemporal dynamics modeling and propose AR-DF for effective training and inference considering intra-frame bidirectionality and inter-frame temporal causality. We anticipate that Lumos-1 represents a significant step toward building a foundational unified model.
Ethics Statement
Our work introduces Lumos-1, an autoregressive model for text-to-video generation. Similar to other generative models, this technology carries societal risks if misused. The primary concerns include the potential for creating deceptive or misleading content for misinformation campaigns, the generation of harmful or disturbing visuals, and the potential amplification of societal biases present in the training data. To address these, we have taken several steps and advocate for further safeguards, as outlined in Sec.˜F.3. We firmly state that Lumos-1 is currently a research-oriented project. We recommend that any future public release of this technology must be preceded by rigorous safety measures, also detailed in Sec.˜F.3.
Reproducibility Statement
We are committed to ensuring the reproducibility of our work. All essential details for reproducing our results are provided within the paper and the appendix. The details of architectures and their training details are presented in and Sec.˜3.3, Sec.˜4.1 and appendix˜D. Our technical contribution, MM-RoPE and AR-DF, are described with comprehensive details in Sec.˜3. The evaluation protocols are detailed in Sec.˜4.1. We will ensure the release of our source code, pre-trained model weights, and evaluation scripts upon publication of this work. We can also provide any source code if any reviewer asks for a detailed implementation.
@@
  • appendix˜A presents more related work and discussions:
    • More related work on autoregressive image generation.
    • The reasons for choosing and retaining the LLM architecture.
    • Discussion on other LLM-based visual generation architectures.
    • Discussion on AR-DF’s relatedness to Diffusion Forcing and FramePack.
  • appendix˜B presents a detailed version introducing vanilla 3D RoPE.
  • appendix˜C presents a detailed version of the AR-DF inference algorithm.
  • appendix˜D presents more architectural and implementation details.
    • Model architectural details and more training details.
    • Detailed introduction of the token sequence formatting.
    • GPU memory friendly implementation.
  • appendix˜E presents more analysis and ablation studies.
    • More ablation studies, including
        The effect of the number of meta MM-RoPEs in MM-RoPE.
        The effect of the scaling factors in MM-RoPE.
    • More analysis, including
        Lumos-1’s robustness to aspect ratios.
        Lumos-1’s comparison with other RoPE designs.
        Lumos-1’s training resource comparison.
    • More visualizations, including
        Qualitative visual comparison on text-to-image generation.
        More text-to-image visualizations.
        More image-to-video and text-to-video visualizations.
  • appendix˜F presents more discussions.
    • A road map to unified models for understanding and generation.
    • Limitations and future work.
    • Potential societal impact and safeguards.
  • appendix˜G presents a disclosure of LLM usage in this paper.
Appendix A More Related Work and Discussions
As a supplement to the main paper, we provide 1) more related work on autoregressive image generation, 2) the reasons for choosing LLMs, 3) the comparison with LLM-based visual generation models and 4) AR-DF’s relatedness to Diffusion Forcing and FramePack.
Autoregressive image generation. Spurred by the progress of LLMs, visual generation research is rapidly shifting from diffusion paradigms [59, 25, 101, 16] to an autoregressive one [65, 71, 94, 29, 56, 104, 98, 24, 47, 79, 57], due to its potential of being integrated into unified models [69, 105] with minimal modification. Preliminary research, such as Parti [95], DALL-E [55], MaskGiT [8] and LLamaGen [65], have demonstrated the efficacy of autoregressive generation using discrete tokenizers. MAR [36] incorporates a diffusion-style objective into autoregressive training, thereby accommodating continuous tokens. VAR [71] and FAR [94] reformulate the learning target as next-scale and next-frequency prediction, respectively, improving generation fidelity. Unlike them, we dive into the design of the RoPE technique in visual generation.
The reasons for choosing and retaining the LLM architecture. One significant group of unified models, like Chameleon [69] and EMU3 [81], adopts the original LLM architecture to unify understanding and generation tasks. Some hybrid architectures combining an autoregressive transformer and a diffusion head, like BLIP3o-Next [12], also rely on the autoregressive visual generation capabilities. Although the autoregressive LLM transformers are very capable language models since they are designed for language tasks, the visual generation results are not satisfactory due to the lack of research exploration. We align our model architecture with the LLM so that the insights obtained and techniques designed (e.g., MM-RoPE and AR-DF) in this paper could transfer smoothly to the unified models to come (but we also want to mention that techniques like MM-RoPE can be transferred to other types of visual generation models). This architectural alignment also means that Lumos-1 is not just a specialized generator; the addition of understanding-oriented data to its pre-training can unlock both generation and comprehension capabilities, which is a promising future work.
Discussion on other LLM-based unified visual generation architectures. We briefly compare our model with representative LLM-based unified generative models, including DALL·E [55], LlamaGen [65], Loong [83] and Lumina-mGPT [43] in Tab.˜A1 to clarify the key distinctions. This comparison highlights three key advantages of Lumos-1’s design: 1) Truly unified architecture: Unlike hybrid models like LlamaGen which stitch together a T5 encoder and an LLM (for visual generation), Lumos-1 uses a single, end-to-end LLM backbone. This is a simpler and more elegant design. 2) Advanced positional encoding (MM-RoPE): Lumos-1 is the only model that incorporates a native 3D-aware RoPE (MM-RoPE) that is also compatible with standard text RoPE used in LLMs. This allows for principled spatiotemporal modeling for visual data, a capability absent in models using 1D, 2D, or non-RoPE embeddings. 3) Efficient prediction paradigm: Crucially, Lumos-1 is the only model in this comparison to leverage fast, parallelizable mask-based prediction. Other models are constrained by the slow, sequential nature of next-token prediction, which becomes prohibitively inefficient for videos. These targeted architectural choices allow Lumos-1 to function as a fast and versatile image-video generator while being consistent with the core LLM structure.
Table A1: Comparison of different LLM-based visual generation models. The table highlights their architecture, position embedding techniques, prediction paradigms, generated media types, and performance on the GenEval benchmark.
Model Architecture Position Embedding Prediction Paradigm Generated Media GenEval
DALL·E LLM $\sin/\cos$ PE Next-token prediction (Slow) Image
LLamaGen T5 + LLM (2.9B + 0.8B) Naïve 2D RoPE Next-token prediction (Slow) Image 0.32
Loong LLM 1D RoPE Next-token prediction (Slow) Image + Video
Lumina-mGPT LLM (7B) 1D RoPE Next-token prediction (Slow) Image 0.56
Lumos-1 LLM (1.5B) MM-RoPE (distributed and scaled 3D RoPE) Mask prediction (Fast) Image + Video 0.725
Discussion on AR-DF’s relatedness to Diffusion Forcing and FramePack. For Diffusion Forcing [9] and AR-DF, they are related since both of them are methods to overcome the fundamental drifting problem in video generation: A small error made early on gets passed down and magnified over time, causing the final video to "drift" far away from the intended or realistic result. The temporal dependency of the video data, to some extent, leads to the drifting problem since later frames’ predictions rely on previous frames. We identify that this is exacerbated by frame-wise loss imbalance since the model tends to optimize those easier tasks, encouraging later frames to rely on "copying" previous frames. The temporal tube masks in AR-DF or the per-frame independent noise in Diffusion Forcing Transformer [63] both aim to ameliorate this issue by reducing the reliance of later frames on previous frames. The inference-time masks in Lumos-1 also mitigate error accumulation by dropping partially generated context. Our temporal tube masking is particularly suited for video data, as it directly targets the spatial information redundancy between frames. This can be revealed by the experiments in Tab.˜5, where we compare our model with a variant trained by following Diffusion Forcing Transformer. We observe that they have similar performance on image generation, while Lumos-1 performs better on video generation. For FramePack [100], it compresses the history in a context with a fixed length. Due to its strategic method for history compression, it focuses more on overcoming the forgetting problem (i.e., later frames forget history context), which is not the focus of Lumos-1. It also proposes anti-drifting sampling to get rid of drifting. However, it relies on access to future frames.
Appendix B Detailed Preliminary of 3D RoPE
One de facto design of contemporary LLM is RoPE [64], whose overall aim is to encode the absolute position with a rotation matrix while incorporating the explicit relative position dependency in the attention mechanism. This can be formulated as:
where $f_{q}(\bm{x}_{m},m)$ encodes the position $m$ for the embedding $\bm{x}_{m}$ to obtain the query feature ($f_{k}(\bm{x}_{n},n)$ is analogously defined); $g(\bm{x}_{m},\bm{x}_{n},m-n)$ is a function that defines the inner product between the query and key vectors that explicitly encodes the relative position mnm-n. The resultant form of function $f_{\{q,k\}}$ can be formulated as:
$f_{\{q,k\}}(\bm{x}_{m},m)=\bm{R}^{d}_{\Theta,m}\bm{W}_{q,k}\bm{x}_{m}$ (9)
$f_{q}(\bm{x}_{m},m)^{\rm T}f_{k}(\bm{x}_{n},n)=\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}\bm{R}^{d}_{\Theta,\tau}\bm{W}_{k}\bm{x}_{n},\quad\tau=n-m$ (10)
$\bm{R}^{d}_{\Theta,\tau}=\begin{bmatrix}R_{\theta_{1},\tau}&0&\cdots&0\\ 0&R_{\theta_{2},\tau}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ 0&0&\cdots&R_{\theta_{d/2},\tau}\\ \end{bmatrix},\quad R_{\theta,\tau}=\begin{bmatrix}\cos\tau\theta&-\sin\tau\theta\\ \sin\tau\theta&\cos\tau\theta\end{bmatrix}$} (11)
However, the application of the original RoPE to modeling visual data remains suboptimal considering the spatiotemporal correlation of visual tokens. One popular type of generative models, diffusion models [26, 25], improved upon this technique by proposing 3D RoPE that jointly injects spatiotemporal latent coordinates during attention calculation and demonstrates its effectiveness [27, 33, 70]. If we slightly abuse the annotation by denoting $\bm{x}_{m}^{\rm T}\bm{W}_{q}^{\rm T}$ and $\bm{W}_{k}\bm{x}_{n}$ as $\bm{X}_{m}^{\rm T}$ and $\bm{X}_{n}$, we can write the attention calculation in equation˜8 based on 3D RoPE as:
$\begin{aligned} \bm{X}_{m,t_{s}:t_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{t_{s}+1},\tau_{t}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{t_{e}},\tau_{t}}\\ \end{bmatrix}\bm{X}_{n,t_{s}:t_{e}}+\bm{X}_{m,h_{s}:h_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{h_{s}+1},\tau_{h}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{h_{e}},\tau_{h}}\\ \end{bmatrix}\bm{X}_{n,h_{s}:h_{e}}+\bm{X}_{m,w_{s}:w_{e}}^{\rm T}\begin{bmatrix}R_{\theta_{w_{s}+1},\tau_{w}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&R_{\theta_{w_{e}},\tau_{w}}\\ \end{bmatrix}\bm{X}_{n,w_{s}:w_{e}}\end{aligned}$
where $\{t_{s},t_{e}\}=\{0,\frac{2}{16}d\}$, $\{h_{s},h_{e}\}=\{\frac{2}{16}d,\frac{5}{16}d\}$ and $\{w_{s},w_{e}\}=\{\frac{5}{16}d,\frac{1}{2}d\}$ denote the start and end dimension index for encoding temporal, height and width relative position; $\bm{X}_{m,t_{s}:t_{e}}^{\rm T}$ denotes the submatrix extracted from $\bm{X}_{m}^{\rm T}$ using row indices $2t_{s},2t_{e})$; other matrices are similarly defined.
Appendix C Detailed AR-DF Inference Algorithm